# A Survey of Low-bit Large Language Models: Basics, Systems, and Algorithms

### v2

@@ 4.1 Low-bit Training
$FP8\_MAX=maximum\_representable\_value(fp8\_format),$ (13)
$exp=get\_exponent(FP8\_MAX/amax),$ (14)
@@
$fp8\_format$ indicates the formats like E4M3 or E5M2. $FP8\_MAX$ is the relevant max value under that format. $amax$ is the maximal absolute value of the tensor. Then we can calculate the $new\_scaling\_factor$ with $exp$. However, the calculation of $new\_scaling\_factor$ can not be online since it will introduce much more memory access. The best practice is to employ delayed scaling. This strategy chooses the scaling factor based on the maximums of absolute values seen in some number of previous iterations. This enables the full performance of FP8 computation but requires storing the history of maximums as additional parameters of the FP8 operators. In Table 3, we list the prevalent frameworks and engines that support low-bit floating-point training, including the Deepspeed [^33] https://github.com/microsoft/DeepSpeed from Microsoft, Megatron-LM [^34] https://github.com/NVIDIA/Megatron-LM from NVIDIA, and UnitScaling [^35] https://github.com/graphcore-research/unit-scaling from GraphCore.
@@
This makes the massive parameter scale of LLMs a more pronounced memory bottleneck during fine-tuning, hindering their deployment in broader application scenarios. INT8 Training (Zhu et al., 2020) is considered a direct method to reduce the memory usage of gradients during training. However, the instability of quantization in backpropagation makes the training of LLMs more unstable and can even lead to crashes. QST (Zhang et al., 2024g) proposes optimizing three key sources of memory usage simultaneously: model weights, optimizer states, and intermediate activations. In addition to quantizing the LLM model weights to 4 bits and introducing a separate side network that uses the LLM’s hidden states for task-specific predictions, QST also uses several low-rank adapters and gradient-free downsampling modules to significantly reduce the number of trainable parameters, thereby saving memory on optimizer states. Q-GaLore (Zhang et al., 2024f) points out that GaLore’s memory-saving strategy of projecting gradients using SVD incurs significant time costs. To address this, Q-GaLore adaptively updates the gradient subspace based on gradient convergence statistics and keeps the projection matrix in INT4 format and the weights in INT8 format, allowing Llama-7b to be trained from scratch on a single 16GB GPU. Jetfire (Xi et al., 2024) features an INT8 data flow to optimize memory access and a per-block quantization method to maintain the accuracy of pretrained transformers. 4-bit Optimizer (Li et al., 2024a) uses a smaller block size and proposes to utilize both row-wise and column-wise information for better quantization, and further identifies a zero point problem of quantizing the second moment, solving it with a linear quantizer.
@@ 4.2 Quantization Strategies for PEFT
The well-pretrained LLMs possess excellent generalization and exhibit good transferability and adaptability during fine-tuning, making them potentially useful for a variety of downstream tasks. However, the massive parameter scale of LLMs creates a significant memory bottleneck during fine-tuning, hindering their broader application. Thus, the concept of parameter-efficient fine-tuning (PEFT) is introduced to address the issue of LLM fine-tuning under resource constraints (Ding et al., 2023; Han et al., 2024).
@@ 4.2.1 Partial Parameter Fine-Tuning with Quantization
Previous QAT methods require almost the same resources as full parameter training, making them infeasible in resource-constrained fine-tuning scenarios. Therefore, partial parameter fine-tuning strategies have been proposed. PEQA (Kim et al., 2024) follows the naive QAT training approach. But after quantizing the weights $W$, it obtains scaling factors $s_{0}$ and fixed-point numbers $\overline{W}_{0}$, then it keeps $W$ and only trains $s_{0}$. OWQ (Lee et al., 2024a) only updates the high-precision “weak columns” after mixed-precision quantization.
@@ 4.2.2 Low-bit Low-Rank Adaptation
Low-Rank Adaptation (LoRA) (Hu et al., 2021) freezes the pre-trained weights and only trains low-rank matrices. Although it reduces the trainable parameters by 10,000 times, it does not decrease the size of the pre-trained model weight itself, thus only reducing the memory requirements for fine-tuning by 3 times.
Methods like QLoRA (Dettmers et al., 2024) utilize low-bit quantization to further reduce the memory occupation by fine-tuning the LoRA for quantized LLMs. They first quantize the pre-trained LLM to low bits using PTQ methods:
$_{q}\leftarrow quant(\textbf{W}),$ W ) , (16)
@@
$\textbf{Y}=\textbf{X}\cdot dequant(\textbf{W})+\textbf{X}\cdot\textbf{A}\textbf{B},$ (17)
@@
In these methods, matrix A is typically initialized with random Gaussian values, while B is initialized to all zeros. This approach not only significantly reduces the memory footprint of the model’s weight parameters but also ensures that the optimizer only needs to store the gradients of LoRA during fine-tuning, greatly decreasing memory usage. QLoRA (Dettmers et al., 2024) introduces the use of Normal Float for double quantization of W, achieving both good accuracy retention and memory savings, allowing fine-tuning of a 65B pre-trained model using a single 48GB GPU. IR-QLoRA (Qin et al., 2024) incorporates information theory into the QLoRA paradigm, enhancing fine-tuning performance through information calibration and connection. LoRA+ (Hayou et al., 2024) demonstrates that setting different learning rates for matrices A and B in LoRA enables efficient feature learning. QDyLoRA (Rajabzadeh et al., 2024) and Bayesian-LoRA (Meo et al., 2024) employs more flexible rank allocation within LoRA.
[…]
Many methods have recognized that the initialization of LoRA significantly impacts the effectiveness of these quantization-based parameter-efficient fine-tuning methods. As a result, they aim to minimize $\left\|\textbf{W}-(\textbf{W}_{q}+\textbf{A}\textbf{B})\right\|_{F}$ before fine-tuning. LoftQ (Li et al., 2023c) and LQ-LoRA (Guo et al., 2023b) both achieve this through iterative computation: $Q_{t}\leftarrow quant(\textbf{W}-\textbf{A}_{t-1}\textbf{B}_{t-1}^{\top})$ and $\textbf{A}_{t},\textbf{B}_{t}\leftarrow SVD(\textbf{W}-Q_{t})$. LQ-LoRA also suggests incorporating calibration data, adjusting the minimization objective to $\left\|\sqrt{F}\odot(\textbf{W}-(\textbf{W}_{q}+\textbf{A}\textbf{B}))\right\|_{F}^{2}$, where $F$ is the Fisher information matrix for W, and $\odot$ represents the Hadamard product. Additionally, LQ-LoRA introduces dynamic quantization configurations to better adapt to resource constraints.
@@
Figure 7 is an illustration of different LoRA structures. Figure 7(a) represents methods like QLoRA that do not alter any part of the LLM during the fine-tuning stage and keep the complete original LoRA structure (Dettmers et al., 2024; Qin et al., 2024; Hayou et al., 2024; Li et al., 2023c). Figure 7(b) represents methods like QA-LoRA that also do not change any part of the LLM during fine-tuning but modify the original LoRA structure (Xu et al., 2023). Figure 7(c) represents methods like L4Q that modify the original LoRA structure and use a training process similar to QAT (Jeon et al., 2024). Both (a) and (b) require only the quantized LLM weights $W_{q}$ during fine-tuning, while (c) needs to store the pre-trained full-precision weights $W_{fp}$. (a) is solely intended to reduce training costs and cannot directly produce a quantized model after fine-tuning, while both (b) and (c) can directly integrate the LoRA module after fine-tuning to produce a deployable quantized model.
@@ 5.1 Quantization-Aware Training
Table 4 summarizes the different QAT methods for LLMs. LLM-QAT (Liu et al., 2023b) is the pioneering work that investigates the QAT for LLMs. To overcome the training data limits, it proposes data-free knowledge distillation which aligns the teacher logits of full-precision models and student logits of quantized models. Following LLM-QAT, BitDistiller (Du et al., 2024) employs the asymmetric clipping strategy for asymmetric quantization during the self-distillation stage. EfficientQAT (Chen et al., 2024b) significantly reduces the training cost by splitting the QAT into two consecutive phases. The first phase optimizes all parameters for each block and then the second phase merely optimizes quantization parameters for the entire network. To pave the way for a new era of extreme quantization level, BitNet (Wang et al., 2023) replaces the BitLinears with original Linears and trains from scratch. Its variant, BitNet b1.58 (Ma et al., 2024a), leverages ternary weight for each parameter which achieves near-lossless performance.
@@ 5.2.1 Equivalent Transformation
Many studies  (Luo et al., 2020; Bondarenko et al., 2021; Wei et al., 2023b; Xiao et al., 2023) have highlighted the presence of significant outliers in LLMs. These outliers pose substantial challenges for quantization, as they force a large number of normal values to be represented with a limited number of bits, which leads to large quantization errors and accuracy degradation. Therefore, a multitude of algorithms have emerged in recent years, aiming to mitigate the issue of outliers in LLMs.
Among all the algorithms addressing the outlier problem, equivalent transformation is one of the most representative and effective methods. One of the pioneering works in applying the equivalent transformation to language models is the Outlier Suppression (OS) (Wei et al., 2022). OS splits the LayerNorm function and migrates $\gamma$, which is a parameter of LayerNorm, to avoid the outlier.
@@
Figure 9: Overall diagram of shifting transformation. $\Delta$ can be merged into the parameter $\beta$ in Layernorm and the weight metrices.
@@
Outliers in LLMs are asymmetrically distributed across different channels. This asymmetrical representation can cause a tensor composed of channels with small ranges to exhibit a very large overall range, resulting in difficulty in the quantization process. To address this issue, OS+ (Wei et al., 2023b) first proposes the channel-wise shifting transformation, which adjusts activations across channels to mitigate the impact of asymmetry as the following equation:
@@
With the channel-wise shifting in place, the tensor range is reduced to the largest channel range, eliminating the influence of asymmetric outliers. However, handcrafting the equivalent parameters leads to sub-optimal results. Hence OmniQuant (Shao et al., 2023) is proposed to determine the optimal shifting parameters in a differentiable way by including the block-wise quantization error minimization:
$\mathop{\arg\min_{\Delta}}||\mathcal{O}(\textbf{W},\textbf{X})-\mathcal{O}\big{(}Q_{w}\left(\textbf{W};\Delta\right),Q_{a}\left(\textbf{X};\Delta\right)\big{)}||,$ (23)
where $\mathcal{O}$ represents the mapping function for a transformer block in the LLM, $Q_{w}(\cdot)$ and $Q_{a}(\cdot)$ denote the weight and activation quantizer respectively, $\Delta$ is the shifting parameter. Block-wise minimization is easy to optimize with minimal resource requirements. Therefore, by optimizing the objective function block by block, a more effective shifting vector can be obtained compared to the direct computation used in OS+ in an efficient and resource-saving way. However, OmniQuant requires fine-tuning of the learnable parameters; otherwise, issues such as gradient explosion can easily occur. Similar to OmniQuant, AffineQuant (Ma et al., 2024b) also adopts a learning-based shifting operation.
@@
Shifting transformation effectively addresses the issue of asymmetrical distribution of outliers in activations, reducing the large range caused by the asymmetry. However, this only aids per-tensor quantization and does not reduce the difficulty of per-channel quantization, as it does not fundamentally eliminate the outliers distributed across channels in the activations. To further reduce the impact of outliers on quantization, SmoothQuant (Xiao et al., 2023) initially proposes to use a scaling transformation. It relies on a key observation: although activations are much more difficult to quantize than weights due to the presence of outliers, different tokens exhibit similar variations across their channels (Dettmers et al., 2022a). Based on this observation, SmoothQuant migrates the quantization difficulty from activations to weights offline by introducing a mathematically equivalent per-channel scaling transformation that significantly smooths the magnitudes across channels:
@@
However, this method requires multiple trials to determine the optimal migration strength for different models, i.e., $\alpha=0.5$ is a well-balanced point for all OPT (Zhang et al., 2022) and BLOOM (Le Scao et al., 2023) models.
Inspired by SmoothQuant, FPTQ (Li et al., 2023b) argues that it is unnecessary to consider weights for computing the activation smoothing scale while it is crucial to retain all the activation values with a non-linear lossless mapping. This mapping needs to fit two criteria: (1) touching gently with the inliers and (2) harshly suppressing the outliers. Based on this, they adopt a logarithmic function to improve the calculation of the smooth matrix $s$:
@@
In addition to FPTQ, many other works have followed the approach of SmoothQuant. Both OS+ and AWQ (Lin et al., 2024a) use searching-based methods to find the smooth scale. However, the optimization objectives and search spaces of the two methods differ. The optimization objective of OS+ is:
$\Phi^{*}=\mathop{\arg\min_{\Phi}}\mathbb{E}||Q\big{(}(\textbf{X}-\Delta)\cdot diag(\Phi)^{-1}\big{)}Q\big{(}diag(\Phi)\cdot\textbf{W}^{\mathsf{T}}\big{)}+\hat{\textbf{b}}-(\textbf{X}\textbf{W}^{\mathsf{T}}+\textbf{b})||^{2}_{F}.$ (27)
To simplify the search space, OS+ optimizes the outlier threshold $t$, compressing channels with an activation range over $t$ into $(-t,t)$ and leaving others unchanged. This reduces the problem to a single variable. A grid search is then used for $t$ to minimize the objective. After finding the optimal $t$, the scaling vector is calculated as follows:
@@
$\Phi={\Phi_{x}}^{\alpha},\quad\alpha^{*}=\mathop{\arg\min_{\alpha}}||Q\big{(}\textbf{W}\cdot diag({\Phi_{x}}^{\alpha})\big{)}(diag({\Phi_{x}}^{\alpha})^{-1}\cdot\textbf{X})-\textbf{W}\textbf{X}||,$ (29)
@@
In addition to searching-based methods, some approaches use learning-based techniques to find the optimal scaling matrix. OmniQuant and AffineQuant also learn the scaling matrix. In Equation 23, OmniQuant learns both the shifting factor $\Delta$ and the scaling matrix $diag(\Phi)$. However, OmniQuant optimizes only within the range of a diagonal matrix. AffineQuant (Ma et al., 2024b) argues that this limited search range can lead to significant quantization errors, reducing the generalizability of the quantization method in low-bit scenarios. It proposes learning a general invertible matrix to perform equivalent affine transformations on weights and activations, achieving better results.
@@
Rotation transformation was first introduced by QuIP (Chee et al., 2024). QuIP is based on the insight that quantization works better when the weight and Hessian matrices are incoherent. This means that the weights should have similar magnitudes and the directions that require precise rounding should not align with the coordinate axes. To make it straight, a weight matrix is $\mu$-incoherent if:
$max(\textbf{W})\leq\mu||\textbf{W}||_{F}/\sqrt{mn},$ (30)
where $mn$ is the number of the matrix elements and $||\cdot||_{F}$ is the Frobenius norm. QuIP shows that multiplying a weight matrix on the left and right by an orthogonal matrix can reduce incoherence, which is equal to performing a rotation transformation on the weight matrix. QuIP utilizes Kronecker-structured orthogonal matrices, allowing for rapid additional computations. Building on this, QuIP# (Tseng et al., 2024a) replaces these with Hadamard matrices, enhancing quantization through better incoherence and speeding up the forward pass, as the Hadamard transform can be computed in $\mathcal{O}(nlogn)$ addition operations.
Both of these two methods target weight-only quantization. Following these approaches, QuaRot (Ashkboos et al., 2024) introduces a weight&activation quantization method that also quantizes the KV cache. QuaRot operates in two stages. First, the model weights are manipulated in full precision, and two Hadamard operations are added to the model’s forward pass. In the second stage, the weights are quantized using an existing method, and quantization operations are integrated into the forward pass for online activation quantization.

### v3

@@ 4.1 Low-bit Training
FP8_MAX=maximum_representable_value(fp8_format),FP8\_MAX=maximum\_representable\_value(fp8\_format), (15)
exp=get_exponent(FP8_MAX/amax),exp=get\_exponent(FP8\_MAX/amax), (16)
@@
fp8_formatfp8\_format indicates the formats like E4M3 or E5M2. FP8_MAXFP8\_MAX is the relevant max value under that format. amaxamax is the maximal absolute value of the tensor. Then we can calculate the new_new\_ scaling_factorscaling\_factor with expexp. However, the calculation of new_scaling_factornew\_scaling\_factor can not be online since it will introduce much more memory access. The best practice is to employ delayed scaling. This strategy chooses the scaling factor based on the maximums of absolute values seen in some number of previous iterations. This enables the full performance of FP8 computation but requires storing the history of maximums as additional parameters of the FP8 operators. Deepseek V3 (deepseekai2024deepseekv3technicalreport), one of the state-of-the-art models, introduces fine-grained block-wise FP8 quantization, enabling highly accurate FP8 training. In Table 3, we list the prevalent frameworks and engines that support low-bit floating-point training, including the Deepspeed [^33] https://github.com/microsoft/DeepSpeed from Microsoft, Megatron-LM [^34] https://github.com/NVIDIA/Megatron-LM from NVIDIA, and UnitScaling [^35] https://github.com/graphcore-research/unit-scaling from GraphCore.
@@
This makes the massive parameter scale of LLMs a more pronounced memory bottleneck during fine-tuning, hindering their deployment in broader application scenarios. INT8 Training (zhu2020towards) is considered a direct method to reduce the memory usage of gradients during training. However, the instability of quantization in backpropagation makes the training of LLMs more unstable and can even lead to crashes. QST (zhang2024quantized) proposes optimizing three key sources of memory usage simultaneously: model weights, optimizer states, and intermediate activations. In addition to quantizing the LLM model weights to 4 bits and introducing a separate side network that uses the LLM’s hidden states for task-specific predictions, QST also uses several low-rank adapters and gradient-free downsampling modules to significantly reduce the number of trainable parameters, thereby saving memory on optimizer states. Q-GaLore (zhang2024q) points out that GaLore’s memory-saving strategy of projecting gradients using SVD incurs significant time costs. To address this, Q-GaLore adaptively updates the gradient subspace based on gradient convergence statistics and keeps the projection matrix in INT4 format and the weights in INT8 format, allowing Llama-7b to be trained from scratch on a single 16GB GPU. Jetfire (xi2024jetfire) features an INT8 data flow to optimize memory access and a per-block quantization method to maintain the accuracy of pretrained transformers. 4-bit Optimizer (li2024memory) uses a smaller block size and proposes to utilize both row-wise and column-wise information for better quantization, and further identifies a zero point problem of quantizing the second moment, solving it with a linear quantizer.
Takeaways of subsection 4.1 BF16 and FP16 training have become widely adopted techniques to accelerate the training process, with relatively lower accuracy risks. FP8, while effective for fine-grained quantization in specific modules like linear layers, carries higher precision risks compared to BF16/FP16. INT8, which has been explored in some research but not yet widely adopted in practice, poses the highest accuracy risks. To mitigate these risks, techniques such as dynamic scaling are often introduced to adjust and stabilize the precision during training.
@@ 4.2 Quantization Strategies for PEFT
The well-pretrained LLMs possess excellent generalization and exhibit good transferability and adaptability during fine-tuning, making them potentially useful for a variety of downstream tasks. However, the massive parameter scale of LLMs creates a significant memory bottleneck during fine-tuning, hindering their broader application. Thus, the concept of parameter-efficient fine-tuning (PEFT) is introduced to address the issue of LLM fine-tuning under resource constraints (ding2023parameter; han2024parameter).
@@ 4.2.1 Partial Parameter Fine-Tuning with Quantization
Previous QAT methods require almost the same resources as full parameter training, making them infeasible in resource-constrained fine-tuning scenarios. Therefore, partial parameter fine-tuning strategies have been proposed. PEQA (kim2024memory) follows the naive QAT training approach. But after quantizing the weights $W$, it obtains scaling factors $s_{0}$ and fixed-point numbers $\overline{W}_{0}$, then it keeps $W$ and only trains $s_{0}$. OWQ (lee2024owq) only updates the high-precision “weak columns" after mixed-precision quantization.
@@ 4.2.2 Low-bit Low-Rank Adaptation
Low-Rank Adaptation (LoRA) (hu2021lora) freezes the pre-trained weights and only trains low-rank matrices. Although it reduces the trainable parameters by 10,000 times, it does not decrease the size of the pre-trained model weight itself, thus only reducing the memory requirements for fine-tuning by 3 times.
Methods like QLoRA (dettmers2024qlora) utilize low-bit quantization to further reduce the memory occupation by fine-tuning the LoRA for quantized LLMs. They first quantize the pre-trained LLM to low bits using PTQ methods:
$\textbf{W}_{q}\leftarrow quant(\textbf{W}),$ (18)
@@
$\textbf{Y}=\textbf{X}\cdot dequant(\textbf{W}_{q})+\textbf{X}\cdot\textbf{A}\textbf{B},$ (19)
@@
In these methods, matrix A is typically initialized with random Gaussian values, while B is initialized to all zeros. This approach not only significantly reduces the memory footprint of the model’s weight parameters but also ensures that the optimizer only needs to store the gradients of LoRA during fine-tuning, greatly decreasing memory usage. QLoRA (dettmers2024qlora) introduces the use of Normal Float for double quantization of W, achieving both good accuracy retention and memory savings, allowing fine-tuning of a 65B pre-trained model using a single 48GB GPU. IR-QLoRA (qin2024accurate) incorporates information theory into the QLoRA paradigm, enhancing fine-tuning performance through information calibration and connection. LoRA+ (hayou2024lora+) demonstrates that setting different learning rates for matrices A and B in LoRA enables efficient feature learning. QDyLoRA (rajabzadeh2024qdylora) and Bayesian-LoRA (meo2024bayesian) employs more flexible rank allocation within LoRA.
Moreover, some methods aim to obtain a deployable quantized and merged model after the LoRA fine-tuning. QA-LoRA (xu2023qa) uses INT format to quantize W and adjusts $\textbf{X}\cdot\textbf{A}^{i\times r}\textbf{B}^{r\times o}$ to $\mathrm{mean}(\textbf{X})\cdot\textbf{A}^{\frac{i}{L}\times r}\textbf{B}^{r\times o}$, allowing the fine-tuned AB to be losslessly merged into the INT format $\textbf{W}_{q}$, without extra computation when deployment. L4Q (jeon2024l4q), on the other hand, maintains the dimension $\textbf{A}\in\mathbb{R}^{i\times r}$ and directly uses the full QAT forward propagation method, simultaneously updating the quantizer parameters $s$ and $b$ for $\textbf{A},\textbf{B}$, and $\textbf{W}+\textbf{A}\textbf{B}$. While L4Q does not reduce the memory footprint of the weights through quantization during pre-training, the optimizer still does not need to retain the gradients of the weights, resulting in a fine-tuned quantized model that can be deployed directly with higher accuracy.
Many methods have recognized that the initialization of LoRA significantly impacts the effectiveness of these quantization-based parameter-efficient fine-tuning methods. As a result, they aim to minimize $\left\|\textbf{W}-(\textbf{W}_{q}+\textbf{A}\textbf{B})\right\|_{F}$ before fine-tuning. LoftQ (li2023loftq) and LQ-LoRA (guo2023lq) both achieve this through iterative computation: $Q_{t}\leftarrow quant(\textbf{W}-\textbf{A}_{t-1}\textbf{B}_{t-1}^{\top})$ and $\textbf{A}_{t},\textbf{B}_{t}\leftarrow SVD(\textbf{W}-Q_{t})$. LQ-LoRA also suggests incorporating calibration data, adjusting the minimization objective to $\left\|\sqrt{F}\odot(\textbf{W}-(\textbf{W}_{q}+\textbf{A}\textbf{B}))\right\|_{F}^{2}$, where $F$ is the Fisher information matrix for W, and $\odot$ represents the Hadamard product. Additionally, LQ-LoRA introduces dynamic quantization configurations to better adapt to resource constraints.
@@
Figure 7 is an illustration of different LoRA structures. Figure 7(a) represents methods like QLoRA that do not alter any part of the LLM during the fine-tuning stage and keep the complete original LoRA structure (dettmers2024qlora; qin2024accurate; hayou2024lora+; li2023loftq). Figure 7(b) represents methods like QA-LoRA that also do not change any part of the LLM during fine-tuning but modify the original LoRA structure (xu2023qa). Figure 7(c) represents methods like L4Q that modify the original LoRA structure and use a training process similar to QAT (jeon2024l4q). Both (a) and (b) require only the quantized LLM weights $W_{q}$ during fine-tuning, while (c) needs to store the pre-trained full-precision weights $W_{fp}$. (a) is solely intended to reduce training costs and cannot directly produce a quantized model after fine-tuning, while both (b) and (c) can directly integrate the LoRA module after fine-tuning to produce a deployable quantized model. Unlike the weight-only quantization in these methods, RoLoRA (huang2024rolora) incorporates rotations with LoRA for effective weight-activation quantization. Although there are existing LoRA works on MoE (li2024mixlora; luo2024moelora; gao2024higher), they have not yet focused on quantization. In the context of quantization, it is crucial to assess whether reducing bit precision exacerbates the expert imbalance problem. Additionally, it is important to explore which position should use the LoRA-MoE method for quantization-aware training (including the router and load balancing) and to examine whether allocating more bits to deeper layers is necessary (gao2024higher).
Takeaways of subsection 4.2 To reduce memory usage during quantization-aware training, a common strategy is to employ partial weight updates, such as updating only a subset of weight columns. To reduce memory usage during normal fine-tuning, quantization can be combined with low-rank approximation techniques, enabling fixed weights to be quantized to lower bit-widths for further memory reduction.
@@ 5.1 Quantization-Aware Training
Table 4 summarizes the different QAT methods for LLMs. LLM-QAT (liu2023llm-qat) is the pioneering work that investigates the QAT for LLMs. To overcome the training data limits, it proposes data-free knowledge distillation which aligns the teacher logits of full-precision models and student logits of quantized models. Following LLM-QAT, BitDistiller (du2024bitdistiller) employs the asymmetric clipping strategy for asymmetric quantization during the self-distillation stage. EfficientQAT (chen2024efficientqat) significantly reduces the training cost by splitting the QAT into two consecutive phases. The first phase optimizes all parameters for each block and then the second phase merely optimizes quantization parameters for the entire network. To pave the way for a new era of extreme quantization level, BitNet (wang2023bitnet) replaces the BitLinears with original Linears and trains from scratch. Its variant, BitNet b1.58 (ma2024era), leverages ternary weight for each parameter which achieves near-lossless performance.
Takeaways of subsection 5.1 Quantization-aware training (QAT) is particularly beneficial in extremely low-bit scenarios, despite its more complex training process. If your goal involves ultra-low-bit configurations and sufficient computational resources are available, QAT can be an effective solution. However, starting QAT from scratch can be challenging; it is generally more practical and efficient to fine-tune a pre-trained model using QAT. Additionally, it is crucial to select training data that generalizes well across diverse domains to mitigate the risk of overfitting.
@@ 5.2.1 Equivalent Transformation
Many studies  (luo2020positional; bondarenko2021understanding; wei2023outlier; xiao2023smoothquant) have highlighted the presence of significant outliers in LLMs. These outliers pose substantial challenges for quantization, as they force a large number of normal values to be represented with a limited number of bits, which leads to large quantization errors and accuracy degradation. Therefore, a multitude of algorithms have emerged in recent years, aiming to mitigate the issue of outliers in LLMs.
Among all the algorithms addressing the outlier problem, equivalent transformation is one of the most representative and effective methods. One of the pioneering works in applying the equivalent transformation to language models is the Outlier Suppression (OS) (wei2022outlier). OS splits the LayerNorm function and migrates $\gamma$, which is a parameter of LayerNorm, to avoid the outlier.
@@
Figure 9: Overall diagram of shifting transformation. $\Delta_{1}$ and $\Delta_{2}$ represent two types of shifting operation. $\Delta_{1}$ can be merged into the parameter $\beta$ in Layernorm and the weight metrics. Specifically, $\Delta_{2}$ can not be merged into the weight matrix. Therefore, the shift transformation between value projection $W_{v}$ and out projection $W_{o}$ can only be conducted online, which may raise extra computation burden.
@@
Outliers in LLMs are asymmetrically distributed across different channels. This asymmetrical representation can cause a tensor composed of channels with small ranges to exhibit a very large overall range, resulting in difficulty in the quantization process. To address this issue, OS+ (wei2023outlier) first proposes the channel-wise shifting transformation, which adjusts activations across channels to mitigate the impact of asymmetry as the following equation:
@@
With the channel-wise shifting in place, the tensor range is reduced to the largest channel range, eliminating the influence of asymmetric outliers. However, handcrafting the equivalent parameters leads to sub-optimal results. Hence OmniQuant (shao2023omniquant) is proposed to determine the optimal shifting parameters in a differentiable way by including the block-wise quantization error minimization:
[…]
where $\mathcal{O}$ represents the mapping function for a transformer block in the LLM, $Q_{w}(\cdot)$ and $Q_{a}(\cdot)$ denote the weight and activation quantizer respectively, $\Delta$ is the shifting parameter. Block-wise minimization is easy to optimize with minimal resource requirements. Therefore, by optimizing the objective function block by block, a more effective shifting vector can be obtained compared to the direct computation used in OS+ in an efficient and resource-saving way. However, OmniQuant requires fine-tuning of the learnable parameters; otherwise, issues such as gradient explosion can easily occur. Similar to OmniQuant, AffineQuant (ma2024affinequant) also adopts a learning-based shifting operation.
@@
Shifting transformation effectively addresses the issue of asymmetrical distribution of outliers in activations, reducing the large range caused by the asymmetry. However, this only aids per-tensor quantization and does not reduce the difficulty of per-channel quantization, as it does not fundamentally eliminate the outliers distributed across channels in the activations. To further reduce the impact of outliers on quantization, SmoothQuant (xiao2023smoothquant) initially proposes to use a scaling transformation. It relies on a key observation: although activations are much more difficult to quantize than weights due to the presence of outliers, different tokens exhibit similar variations across their channels (dettmers2022gpt3). Based on this observation, SmoothQuant migrates the quantization difficulty from activations to weights offline by introducing a mathematically equivalent per-channel scaling transformation that significantly smooths the magnitudes across channels:
@@
However, this method requires multiple trials to determine the optimal migration strength for different models, i.e., $\alpha=0.5$ is a well-balanced point for all OPT (zhang2022opt) and BLOOM (le2023bloom) models.
Inspired by SmoothQuant, FPTQ (li2023fptq) argues that it is unnecessary to consider weights for computing the activation smoothing scale while it is crucial to retain all the activation values with a non-linear lossless mapping. This mapping needs to fit two criteria: (1) touching gently with the inliers and (2) harshly suppressing the outliers. Based on this, they adopt a logarithmic function to improve the calculation of the smooth matrix $s$:
@@
In addition to FPTQ, many other works have followed the approach of SmoothQuant. Both OS+ and AWQ (lin2024awq) use searching-based methods to find the smooth scale. However, the optimization objectives and search spaces of the two methods differ. The optimization objective of OS+ is:
$\Phi^{*}$ $=$ $\mathop{\arg\min_{\Phi}}\mathbb{E}||Q\big((\textbf{X}-\Delta)\cdot diag{(\Phi)}^{-1}\big)Q\big(diag(\Phi)\cdot\textbf{W}^{\mathsf{T}}\big)$ (29)
$+\hat{\textbf{b}}-(\textbf{X}\textbf{W}^{\mathsf{T}}+\textbf{b})||^{2}_{F}.$
To simplify the search space, OS+ optimizes the outlier threshold $t$, compressing channels with an activation range over $t$ into (t,t)(-t,t) and leaving others unchanged. This reduces the problem to a single variable. A grid search is then used for $t$ to minimize the objective. After finding the optimal $t$, the scaling vector is calculated as follows:
@@
$\Phi$ $={\Phi_{x}}^{\alpha},$
$\alpha^{*}$ $=\mathop{\arg\min_{\alpha}}\left\|Q\!\left(\textbf{W}\cdot\mathrm{diag}({\Phi_{x}}^{\alpha})\right)(\mathrm{diag}({\Phi_{x}}^{\alpha}))^{-1}\textbf{X}-\textbf{W}\textbf{X}\right\|,$ (31)
@@
In addition to searching-based methods, some approaches use learning-based techniques to find the optimal scaling matrix. OmniQuant and AffineQuant also learn the scaling matrix. In Equation 25, OmniQuant learns both the shifting factor $\Delta$ and the scaling matrix $diag(\Phi)$. However, OmniQuant optimizes only within the range of a diagonal matrix. AffineQuant (ma2024affinequant) argues that this limited search range can lead to significant quantization errors, reducing the generalizability of the quantization method in low-bit scenarios. It proposes learning a general invertible matrix to perform equivalent affine transformations on weights and activations, achieving better results.
@@
Rotation transformation was first introduced by QuIP (chee2024quip). QuIP is based on the insight that quantization works better when the weight and Hessian matrices are incoherent. This means that the weights should have similar magnitudes and the directions that require precise rounding should not align with the coordinate axes. To make it straight, a weight matrix is $\mu$-incoherent if:
[…]
where $mn$ is the number of the matrix elements and $||\cdot||_{F}$ is the Frobenius norm. QuIP shows that multiplying a weight matrix on the left and right by an orthogonal matrix can reduce incoherence, which is equal to performing a rotation transformation on the weight matrix. QuIP utilizes Kronecker-structured orthogonal matrices, allowing for rapid additional computations. Building on this, QuIP# (tseng2024quip) replaces these with Hadamard matrices, enhancing quantization through better incoherence and speeding up the forward pass, as the Hadamard transform can be computed in $\mathcal{O}(nlogn)$ addition operations.
Both of these two methods target weight-only quantization. Following these approaches, QuaRot (ashkboos2024quarot) introduces a weight&activation quantization method that also quantizes the KV cache. QuaRot operates in two stages. First, the model weights are manipulated in full precision, and two Hadamard operations are added to the model’s forward pass. In the second stage, the weights are quantized using an existing method, and quantization operations are integrated into the forward pass for online activation quantization.